# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 1.5. Vistas
The mod $\ell$ non-vanishing of self-dual Hecke $L$-values has further applications. For example, it is ancillary to the main result of [37] on primitivity of Hida family of theta lifts from the unitary group $U(1)$ to a totally definite $U(2)$ over a CM field $K$. It is also a key ingredient in the joint work [1] of the first-named author with Alonso Rodrigues and Lai on primitivity of the Eisenstein series ${\mathcal{E}}(\psi)$ on $U(2,1)$ over $K$ by a new method. The loc. cit. considers periods of ${\mathcal{E}}(\psi)$ against the subgroup $U(1)\times U(2)\subset U(2,1)$ instead of Fourier–Jacobi coefficients as in [28] (cf. §1.3.1). Such a period against a CM form on $U(2)$ is a product of Hecke $L$-values. This approach may lead to removal of some of the hypotheses from the CM main conjecture divisibility (cf. Theorem 1.6).
The rigidity of the zero divisor of Katz $p$-adic $L$-function as in Theorem 1.9 is perhaps an example of a general phenomenon. In the sequel [5] we prove an analogous rigidity for anticyclotomic Rankin–Selberg $p$-adic $L$-functions over a CM field.
A basic problem: mod $\ell$ non-vanishing of Hecke $L$-values over a CM field for non-ordinary primes $\ell$. Note that our mod $\ell$ non-vanishing of Rankin–Selberg $L$-values as in Theorem 1.11 allows $\ell$ to be non-ordinary. In conjunction with Finis’ ideas [14] on the arithmetic of theta functions with complex multiplication, we plan to study this problem.
@@ 1.6. Plan of the paper
In section 2 we recall generalities regarding modular forms and special points on a Shimura set over a totally real field. Then section 3 describes an explicit Waldspurger formula connecting Rankin–Selberg central $L$-values in an anticyclotomic family and toric periods on the Shimura set, leading to an $\ell$-integrality of these $L$-values. Sections 4 and 5 constitute the core of the paper. In section 4 we prove mod $\ell$ non-vanishing of the Rankin–Selberg central $L$-values in the CM case, via refined Fourier analysis and Ratner’s ergodicity of unipotent flows. Then section 5 establishes the main result on mod $\ell$ non-vanishing of self-dual Hecke $L$-values. In the last section we describe an application to the CM Iwasawa main conjecture.
@@ 1.7. Notation
  • Throughout the paper, we fix totally real number field $F$. Let $D_{F}$, $O$ and $\mathbb{A}$} respectively denote the absolute discriminant, the ring of integers and the ring of adéles of $F$. Let ${\mathbb{A}}_{f}$ denote the ring of finite adéles of $F$.
  • We also fix a totally imaginary quadratic extension $K/F$.
@@
  • For a finite place $v$ of $F$, let $F_{v}$ be the completion of $F$ at $v$. Let $O_{v}$ denote the ring of integers of $F_{v}$, $\varpi_{v}$ a uniformiser of $O_{v}$ and ${\mathfrak{p}}_{v}$ the prime ideal of $O_{v}$. Let $q_{v}$ be the cardinality of residue field of $O_{v}$. We also fix a generator $d_{v}$ of the different $D_{v}$ of $F_{v}$ and $d_{K_{v}/F_{v}}$ a generator of the relative different $D_{K_{v}/F_{v}}$ of $K_{v}/F_{v}$. For an ideal $\mathfrak{c}$} of $F$, let $O_{K,{\mathfrak{c}}}=O_{F}+{\mathfrak{c}}O_{K}$ be the $O_{F}$-order of conductor $\mathfrak{c}$}.
@@
  • For a finite place $v$ of $F$ and $\chi_{v}$ a character of $K_{v}$ that is trivial on $O_{v}^{\times}$, define its conductor $\mathrm{cond}(\chi_{v})$ with respect to $F_{v}$ as the maximal integral ideal ${\mathfrak{c}}_{v}$ of $F_{v}$ such that
    $\chi_{v}(O_{K_{v},{\mathfrak{c}}_{v}}^{\times})=1.$
    For a Hecke character $\chi$ over $K$ trivial on $\mathbb{A}$$^{\times}$, the global conductor $\mathrm{cond}(\chi)$ is analogously defined. We let ${\mathrm{Cond}}(\lambda)$ denote the usual notion of conductor of a Hecke character $\lambda$ over $K$, which is the maximal integral ideal $\mathfrak{f}$} of $K$ such that $\lambda((1+\widehat{{\mathfrak{f}}})^{\times})=1$.
@@ Acknowledgments
We thank Haruzo Hida and Christopher Skinner for numerous instructive discussions. We also thank Raul Alonso Rodrigues, Shinichi Kobayashi, Shilin Lai, Lei Yang, Shou-Wu Zhang and Wei Zhang for helpful discussions.
The work of A.B. is partially supported by the NSF grant DMS 2302064, W.H. by the NSFC grant No. 12501019 and Y.T. by the NSFC grant No. 12288201.
@@ 2.1.1. CM field
2.1.2. Totally definite quaternion algebra
@@
Suppose that
@@ 2.1.3. Algebraic representations of $G(F)$
Fix an isomorphism
@@
For any subalgebra $A$ of $\overline{{\mathbb{Q}}}$ that contains the Galois closure of $F$ so that $\iota$ as in (2.4) is defined over $A$, note that $\rho_{k}$ is also defined over $A$. Put
@@ 2.1.5. $\ell$-adic avatar of modular forms
Let $A\subset\overline{{\mathbb{Q}}}_{\ell}$ be a complete algebra that contains $\iota_{\ell}\circ\sigma(O)$ for all $\sigma\in\Sigma$ and so that $\iota$ as in (2.4) is defined over ${\mathrm{Frac}}(A)\cap\overline{{\mathbb{Q}}}$. Then we have a representation $(\rho_{k},V_{k-2\Sigma}(A))$ of $\rho(F)$ defined similarly as in §2.1.3. Let $U=\prod_{v}U_{v}\subset\widehat{B}^{\times}$ be an open compact subgroup such that $\rho_{k,\ell}|_{\prod_{v|\ell}{U_{v}}}$ preserves $V_{k-2\Sigma}(A)$.
@@ 2.1.6. $\ell$-optimal form
$\langle{x,\widehat{\varphi}(g)}\rangle_{x}\in O_{\tilde{L}}$
for all $g\in\widehat{B}^{\times}$, and there exists a $g\in\widehat{B}^{\times}$ such that $\langle{x,\widehat{\varphi}(g)}\rangle_{x}$ is an $\ell$-adic unit.
@@
If $\ell\Sigma>(k-2)\in{\mathbb{Z}}[\Sigma]$, then $\widehat{\varphi}$ is $\ell$-optimal with respect to $x$ if and only if $\widehat{\varphi}$ is $\ell$-integral and is non-zero modulo $\ell$.
@@ 2.2.1. Toric embedding
$\varsigma_{v}=$ $\delta^{-1}\begin{pmatrix}\theta&c(\theta)\\ 1&1\\ \end{pmatrix}\in{\mathrm{GL}}_{2}(K_{w})={\mathrm{GL}}_{2}(F_{v})\text{ if $v\in S_{1}^{+}\backslash S_{0}$.}$
For each $v\in S_{1}^{+}$, choose a place $w$ of $K$ above $v$. If $v\in S_{1}^{+}\backslash S_{0}$, for $t=(t_{1},t_{2})\in K_{v}:=K\otimes_{F}F_{v}=K_{w}\oplus K_{\overline{w}}$ we have
@@
For each $v\in S_{0}$ and a non-negative integer $c_{v}$, define $\varsigma_{v}^{(c_{v})}\in G(F_{v})$ as follows.
Fix an integer $m$, let $e$ be the ramification index of $F_{v}/{\mathbb{Q}}_{p}$. For $n\geq m$, let $0\leq j\leq e-1$ be an integer such that $m+j\equiv n\pmod{e}$ and put $\delta_{n}=\begin{pmatrix}\varpi_{v}^{m+j}p^{\frac{n-m-j}{e}}&\\ &1\end{pmatrix}$.
@@
$\varsigma_{v}^{(c_{v})}:=$ $\begin{pmatrix}\theta&-1\\ 1&0\\ \end{pmatrix}\delta_{c_{v}}\in{\mathrm{GL}}_{2}(K_{\wp})={\mathrm{GL}}_{2}(F_{v}).$ (2.7)
@@
$\varsigma_{v}^{(c_{v})}:=$ $\begin{pmatrix}0&1\\ -1&0\\ \end{pmatrix}\delta_{c_{v}}.$ (2.8)
We take $\varsigma_{\infty}=1$ and for ${\mathfrak{c}}=\prod_{v\in S_{0}}{\mathfrak{p}}_{v}^{c_{v}}$, let $\varsigma^{({\mathfrak{c}})}=\prod_{v\notin S_{0}}\varsigma_{v}\prod_{v\in S_{0}}\varsigma_{v}^{({\mathrm{ord}}_{v}{\mathfrak{c}})}$.
@@
$\iota_{\varsigma^{({\mathfrak{c}})}}:{\mathbb{A}}_{K}\hookrightarrow B({\mathbb{A}}).$
@@ 2.2.2. Quaternionic order
Let ${\mathfrak{c}}=\prod_{v\in S_{0}}{\mathfrak{p}}_{v}^{c_{v}}$ such that $c_{v}\geq n_{v}$ for any $v\in S_{0}$.
Together with the choice of $\varsigma^{({\mathfrak{c}})}$ as above, we have
$\widehat{R}\cap\iota_{\varsigma^{({\mathfrak{c}})}}\widehat{K}=\iota_{\varsigma^{({\mathfrak{c}})}}\widehat{O}_{K,{\mathfrak{c}}}.$
@@ 2.2.3. Special points
Define $x_{{\mathfrak{c}}}:{\mathbb{A}}_{K}^{\times}\to G({\mathbb{A}})$ by
$x_{{\mathfrak{c}}}(a):=a\cdot\varsigma^{({\mathfrak{c}})}\quad(\varsigma^{({\mathfrak{c}})}:=\prod_{v\notin S_{0}}\varsigma_{v}\prod_{v\in S_{0}}\varsigma_{v}^{({\mathrm{ord}}_{v}{\mathfrak{c}})}\varsigma_{v}).$ (2.9)
This gives a family of special points $\{x_{{\mathfrak{c}}}(a)\}_{a\in{\mathbb{A}}_{K}^{\times}}$.
@@ 3.1.1. Setting
Here $\varepsilon(\sigma_{K})$ is the epsilon factor [^8] See [36, Chap. 2] for base change representations and [33, Thm. 2.18] for definition of epsilon factors of ${\mathrm{GL}}_{2}$ representations. of base change $\sigma_{K}$.
@@
For an ideal $\mathfrak{c}$} of $F$, put
${\rm G}_{\mathfrak{c}}=K^{\times}\widehat{F}^{\times}\backslash\widehat{K}^{\times}/\widehat{O}_{K,{\mathfrak{c}}}^{\times}.$
@@
where $\mathfrak{c}$} runs over all ideals of $F$ that are supported on $S_{0}$. Let ${\mathcal{X}}_{S_{0}}$ be the set of $\overline{{\mathbb{Q}}}^{\times}$-valued finite order characters $\chi$ of ${\rm G}_{S_{0}}$ such that
@@ 3.1.2. Some consequences of the Tunnell–Saito theorem
If $v\in S_{0}$, then the non-vanishing of ${\mathrm{Hom}}_{K_{v}^{\times}}(\pi_{v},\chi_{v})$ is a consequence of the explicit Tunnell–Saito theorem as in  [9, Prop. 3.1].
@@
Suppose that $\pi$ has CM by $K$ in the sense of Example 1. Then the quaternion algebra $B$ as in (3.1) satisfies
@@
Suppose that $\sigma=\sigma_{\lambda}$ is associated to $\lambda$ as in Example 1. For each $v$, let $\psi_{v}$ be a non-trivial additive character of $F_{v}$ and $\psi_{K_{v}}=\psi_{v}\circ{\mathrm{tr}}_{K_{v}/F_{v}}$. Let $\varepsilon(\sigma_{v},\psi_{v})$ be the epsilon factor of $\sigma_{v}$ with respect to $\psi_{v}$ defined in [33, Thm. 2.18] and $\varepsilon(\sigma_{K_{v}},\psi_{K_{v}})$ the base change epsilon factor.
@@
Here $\lambda_{v}^{*}:=\lambda_{v}\cdot|\cdot|_{K_{v}}^{1/2}$ denotes the unitarisation of $\lambda_{v}$, the second equality follows from [33, Thm. 4.7] and the last from
@@
By the lemma, the totally definite quaternion algebra $B$ is an intrinsic invariant of the CM field $K$.

### v2

@@ 1.5. Vistas
The mod $\ell$ non-vanishing of self-dual Hecke $L$-values has various applications. For example, it is ancillary to the main result of [35] on primitivity of Hida family of theta lifts from the unitary group $U(1)$ to a totally definite $U(2)$ over a CM field $K$. It is also a key ingredient in the joint work [1] of the first-named author with Alonso Rodrigues and Lai on primitivity of the Eisenstein series ${\mathcal{E}}(\psi)$ on the unitary group $U(2,1)$ over $K$ by a new method. The loc. cit. considers periods of ${\mathcal{E}}(\psi)$ against the subgroup $U(1)\times U(2)\subset U(2,1)$ instead of Fourier–Jacobi coefficients as in [27] (cf. §1.3.1). Such a period against a CM form on $U(2)$ is a product of Hecke $L$-values. This approach may lead to removal of some of the hypotheses from the CM main conjecture divisibility (cf. Theorem 1.6).
In a sequel [4] we will approach the non-vanishing problems (Q1) and (Q2) when the prime $\mathfrak{p}$} is of degree greater than one via a refinement of the current strategy [^8] The current strategy readily yields the non-vanishing for a Zariski dense subset of characters.. The connection between epsilon factor of a self-dual Hecke character and non-vanishing of associated test vector on a specific component of the Shimura set established in §4.4.2 can be viewed as a global refinement of the local Tunnell–Saito theorem. It suggests a new structure on the endoscopic cases of the global Gan–Gross–Prasad conjecture, which we will also explore.
A basic problem: mod $\ell$ non-vanishing of Hecke $L$-values over a CM field for non-ordinary primes $\ell$. Note that our mod $\ell$ non-vanishing of Rankin–Selberg Hecke $L$-values as in Theorem 1.9 allows $\ell$ to be non-ordinary. In conjunction with Finis’ ideas [13] on the arithmetic of theta functions with complex multiplication, we plan to study this problem.
@@ 1.6. Plan of the paper
In section 2 we recall generalities regarding modular forms and special points on a Shimura set over a totally real field. Then section 3 describes an explicit Waldspurger formula connecting Rankin–Selberg central $L$-values in an anticyclotomic family and toric periods on the Shimura set, leading to an $\ell$-integrality of these $L$-values. Sections 4 and 5 constitute the core of the paper. In section 4 we prove mod $\ell$ non-vanishing of the Rankin–Selberg central $L$-values in the CM case, via automorphic representation theory and Ratner’s ergodicity of unipotent flows. Section 5 establishes the main result on mod $\ell$ non-vanishing of self-dual Hecke $L$-values. In the last section we describe an application to the CM Iwasawa main conjecture.
@@ 1.7. Notation
  • Throughout the paper, we fix totally real number field $F$. Let $D_{F}$, $O$ and $\mathbb{A}$} respectively denote the absolute discriminant, the ring of integers and the ring of adéles of $F$. Let ${\mathbb{A}}_{f}$ denote the ring of finite adéles of $F$. For a finite set of places $S$ of $F$, let ${\mathbb{A}}^{(S)}\subset{\mathbb{A}}$ denote the adele ring with the factors at $S$ omitted.
  • We also fix a totally imaginary quadratic extension $K/F$, and analogously define $D_{K}$, $O_{K}$ and ${\mathbb{A}}_{K}$. Let $D_{K/F}$ denote the relative discriminant of $K/F$.
@@
  • For a finite place $v$ of $F$, let $F_{v}$ be the completion of $F$ at $v$. Let $O_{v}$ denote the ring of integers of $F_{v}$, $\varpi_{v}$ a uniformiser of $O_{v}$ and ${\mathfrak{p}}_{v}$ the prime ideal of $O_{v}$. Let $q_{v}$ be the cardinality of residue field of $O_{v}$. We also fix a generator $d_{v}$ of the different $D_{v}$ of $F_{v}$ and $d_{K_{v}/F_{v}}$ a generator of the relative different $D_{K_{v}/F_{v}}$ of $K_{v}/F_{v}$. For an ideal $\mathfrak{c}$} of $F$, let $O_{K,{\mathfrak{c}}}=O+{\mathfrak{c}}O_{K}$ be the $O$-order of conductor $\mathfrak{c}$}.
@@
  • For a finite place $v$ of $F$ and $\chi_{v}$ a character of $K_{v}$, define its conductor $\mathrm{cond}(\chi_{v})$ with respect to $F_{v}$ as the maximal integral ideal ${\mathfrak{r}}_{v}$ of $F_{v}$ such that
    $\chi_{v}((1+{\mathfrak{r}}_{v}O_{K_{v}})^{\times})=1.$
    For a Hecke character $\chi$ over $K$, the global conductor $\mathrm{cond}(\chi)$ is analogously defined. We let ${\mathrm{Cond}}(\chi)$ denote the usual notion of conductor of a Hecke character $\chi$ over $K$, which is the maximal integral ideal $\mathfrak{f}$} of $K$ such that $\chi((1+\widehat{{\mathfrak{f}}})^{\times})=1$.
@@ Acknowledgments
We thank Haruzo Hida and Christopher Skinner for numerous instructive discussions. We also thank Shilin Lai and Wei Zhang for helpful comments on the preprint, and Raul Alonso Rodrigues, Li Cai, Shinichi Kobayashi, Philippe Michel and Lei Yang for discussions.
The work of A.B. is partially supported by the NSF grant DMS 2302064, W.H. by the NSFC grant No. 12501019, Y.T. by the NSFC grant No. 12288201 and X.Y. by NSFC grant No. 12031019.
@@ 2.1.1. CM field
2.2. Totally definite quaternion algebra
@@
Suppose that [^9] The hypotheses will be satisfied in our setting (cf. Lemma 3.2).
@@
2.3. Definite quaternionic modular forms
In this subsection we briefly describe algebraic and $\ell$-integral theory of modular forms on the multiplicative group of a totally definite quaternion algebra.
@@ 2.3.1. Algebraic representations of $G(F)$
Let $B$ be a totally definite quaternion algebra over $F$. Fix an isomorphism
@@
For any subalgebra $A$ of $\overline{{\mathbb{Q}}}$ that contains the Galois closure of $F$ so that $\iota$ is defined over $A$, note that $\rho_{k}$ is also defined over $A$. Put
@@
We may and do choose $\iota$ such that (2.4) holds and $x_{0}$ is $K^{\times}$-invariant.
@@ 2.3.3. $\ell$-adic avatar of modular forms
Let $A\subset\overline{{\mathbb{Q}}}_{\ell}$ be a complete algebra that contains $\iota_{\ell}\circ\sigma(O)$ for all $\sigma\in\Sigma$ and so that $\iota$ as in (2.4) is defined over ${\mathrm{Frac}}(A)\cap\overline{{\mathbb{Q}}}$. Then we have a representation $(\rho_{k},V_{k-2\Sigma}(A))$ of $\rho(F)$ defined similarly as in §2.3.1. Let $U=\prod_{v<\infty}U_{v}\subset\widehat{B}^{\times}$ be an open compact subgroup such that $\rho_{k,\ell}|_{\prod_{v|\ell}{U_{v}}}$ preserves $V_{k-2\Sigma}(A)$.
@@ 2.3.4. $\ell$-optimal form
$\langle{x,\widehat{\varphi}(g)}\rangle_{k-2\Sigma}\in O_{\tilde{L}}$
for all $g\in\widehat{B}^{\times}$, and there exists a $g\in\widehat{B}^{\times}$ such that $\langle{x,\widehat{\varphi}(g)}\rangle_{k-2\Sigma}$ is an $\ell$-adic unit.
@@
If $\ell\Sigma>(k-2\Sigma)\in{\mathbb{Z}}[\Sigma]$, then $\widehat{\varphi}$ is $\ell$-optimal with respect to $x$ if and only if $\widehat{\varphi}$ is $\ell$-integral and is non-zero [^10] Since $G({\mathbb{F}}_{\ell^{r}})$ representation $V_{k_{\sigma}-2}({\mathbb{F}}_{\ell^{r}})$ is irreducible for $r\in{\mathbb{Z}}_{\geq 0}$ and $\sigma|\infty$. modulo $\ell$.
@@ 2.4.1. Toric embedding
$\varsigma_{v}=$ $\delta^{-1}\begin{pmatrix}\theta&c(\theta)\\ 1&1\\ \end{pmatrix}\in{{\mathrm{SL}}_{2}(O_{K_{w}})}={{\mathrm{SL}}_{2}(O_{F_{v}})}\text{ if $v\in S_{1}^{+}\backslash S_{0}$,}$
where $\theta$ and $\delta$ are as in §2.1.1. For each $v\in S_{1}^{+}$, choose a place $w$ of $K$ above $v$. If $v\in S_{1}^{+}\backslash S_{0}$, for $t=(t_{1},t_{2})\in K_{v}:=K\otimes_{F}F_{v}=K_{w}\oplus K_{\overline{w}}$ we have
@@
For each $v\in S_{0}$ and a non-negative integer $r_{v}$, define $\varsigma_{v}^{(r_{v})}\in G(F_{v})$ as follows.
To begin, for $r_{v}\geq 0$, put $\delta_{r_{v}}=\begin{pmatrix}\varpi_{v}^{r_{v}}&\\ &1\end{pmatrix}$.
@@
$\varsigma_{v}^{(r_{v})}:=$ $\begin{pmatrix}\theta&-1\\ 1&0\\ \end{pmatrix}\delta_{r_{v}}\in{\mathrm{GL}}_{2}(K_{\wp})={\mathrm{GL}}_{2}(F_{v}).$ (2.7)
@@
$\varsigma_{v}^{(r_{v})}:=$ $\begin{pmatrix}0&1\\ -1&0\\ \end{pmatrix}\delta_{r_{v}}.$ (2.8)
We take $\varsigma_{\infty}=1$ and for ${\mathfrak{r}}=\prod_{v\in S_{0}}{\mathfrak{p}}_{v}^{r_{v}}$, let $\varsigma^{({\mathfrak{r}})}=\prod_{v\notin S_{0}}\varsigma_{v}\prod_{v\in S_{0}}\varsigma_{v}^{(r_{v})}$.
@@
$\iota_{\varsigma^{({\mathfrak{r}})}}:{\mathbb{A}}_{K}\hookrightarrow B({\mathbb{A}}).$
The above explicit embeddings satisfy (2.9) and Lemma 4.22, which will be used in our non-vanishing arguments.
@@ 2.4.2. Quaternionic order
Let ${\mathfrak{r}}=\prod_{v\in S_{0}}{\mathfrak{p}}_{v}^{r_{v}}$ such that $r_{v}\geq n_{v}$ for any $v\in S_{0}$.
Together with the choice of $\varsigma^{({\mathfrak{r}})}$ as above, we have
$\widehat{R}\cap\iota_{\varsigma^{({\mathfrak{r}})}}\widehat{K}=\iota_{\varsigma^{({\mathfrak{r}})}}\widehat{O}_{K,{\mathfrak{r}}}.$ (2.9)
@@ 2.4.3. Special points
Define $x_{{\mathfrak{r}}}:{\mathbb{A}}_{K}^{\times}\to G({\mathbb{A}})$ by
$x_{{\mathfrak{r}}}(a):=a\cdot\varsigma^{({\mathfrak{r}})},\quad\varsigma^{({\mathfrak{r}})}:=\prod_{v\notin S_{0}}\varsigma_{v}\prod_{v\in S_{0}}\varsigma_{v}^{({\mathrm{ord}}_{v}{\mathfrak{r}})}\varsigma_{v}.$ (2.10)
This gives a family of special points $\{x_{{\mathfrak{r}}}(a)\}_{a\in{\mathbb{A}}_{K}^{\times}}$.
@@ 3.1.1. Setting
Here $\varepsilon(\sigma_{K})$ is the epsilon factor [^11] See [34, Chap. 2] for base change representations and [31, Thm. 2.18] for definition of epsilon factors of ${\mathrm{GL}}_{2}$ representations. of base change $\sigma_{K}$.
@@
For an ideal $\mathfrak{r}$} of $F$, put
${\rm P}_{\mathfrak{r}}=K^{\times}\widehat{F}^{\times}\backslash\widehat{K}^{\times}/\widehat{O}_{K,{\mathfrak{r}}}^{\times}.$
@@
where $\mathfrak{r}$} runs over all ideals of $F$ that are supported on $S_{0}$. Let ${\mathcal{X}}_{S_{0}}$ be the set of $\overline{{\mathbb{Q}}}^{\times}$-valued finite order characters $\chi$ of ${\rm P}_{S_{0}}$ such that
@@ 3.1.2. Some consequences of the Tunnell–Saito theorem
If $v\in S_{0}$, then the non-vanishing of ${\mathrm{Hom}}_{K_{v}^{\times}}(\pi_{v},\chi_{v})$ is a consequence of the explicit Tunnell–Saito theorem as in  [8, Prop. 3.1].
@@
Suppose that $\sigma$ has CM by $K$ in the sense of Example 1. Then the quaternion algebra $B$ as in (3.1) satisfies [^13] In particular, the hypothesis of subsection 2.2 is satisfied.
@@
Suppose that $\sigma=\sigma_{\lambda}$ is associated to $\lambda$ as in Example 1. For each $v$, let $\psi_{v}$ be a non-trivial additive character of $F_{v}$ and $\psi_{K_{v}}=\psi_{v}\circ{\mathrm{tr}}_{K_{v}/F_{v}}$. Let $\varepsilon(\sigma_{v},\psi_{v})$ be the epsilon factor of $\sigma_{v}$ with respect to $\psi_{v}$ defined in [31, Thm. 2.18] and $\varepsilon(\sigma_{K_{v}},\psi_{K_{v}})$ the base change epsilon factor.
@@
Here $\lambda_{v}^{*}:=\lambda_{v}\cdot|\cdot|_{K_{v}}^{1/2}$ denotes the unitarisation of $\lambda_{v}$, the second equality follows from [31, Thm. 4.7] and the last from
@@
By Lemma 3.2, the totally definite quaternion algebra $B$ is an intrinsic invariant of the CM field $K$.